# MetaLint: Generalizable Idiomatic Code Quality Analysis Through Instruction-Following and Easy-to-Hard Generalization

### v2

@@ 3.2 Instruction Supervised Fine-Tuning
As discussed in Section 3.1, we train the target LLM $\Phi$ on a set of linter-detectable, easy idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$. The input to the model consists of a prompt $p$, which combines a meta-task specification $M_{I}$ for some $I\in I_{\mathcal{L}}$ with a source code file $f\in\mathcal{F}$. The model's output is a list of idiom violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix C.1). In cases where there are no violations ($|V_{f,I}|=0$), the model is expected to output the phrase NO VIOLATIONS FOUND. We attempt to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some Python idioms, the final distribution is approximately 70:30 in favor of files with no violations for Python data, but roughly 53:47 (PMD) and 50:50 (JEP Tree-Sitter) for the Java data. This results in a total of 53k synthetic training instances spanning 50 idioms (a subset of all the idioms detectable by Ruff) for Python Ruff data and 96.8k instances spanning 269 idioms for Java PMD and 127.3k instances spanning 15 idioms for tree-sitter data, respectively.
@@ 3.3 Verifiable Reward Model and Preference Optimization
For preference optimization, we adopt the RS-DPO approach (Khaki et al., 2024), which combines rejection sampling (RS) (Touvron et al., 2023) with Direct Preference Optimization (DPO) (Rafailov et al., 2023) to generate on-policy data from a supervised fine-tuned (SFT) policy model. It samples $k$ outputs per input, computes rewards for them, and constructs contrastive win–loss pairs based on the reward distribution and a threshold $\eta$ (Figure 2). We detail the verifiable linter-based reward model and contrastive pair sampling procedure below.
Reward Model Design: The reward model evaluates model outputs by comparing predicted violations against those flagged by the linter, treating the linter's line numbers (blue circle in ``Verifiable Reward Model'', Figure 1) as ground truth and the model’s predicted lines (yellow circle) as predictions. Reward is computed using set-based precision, recall, and F1-score (visualized via the Venn diagram in the same figure), based on line-level overlap. Since each meta-task $M_{I}$ corresponds to a single idiom $I$, we compute one F1-score (reward) per instance.
Sampling Contrastive Pairs: We begin with an SFT policy model $\Phi^{SFT}$ and sample $k=5$ outputs $y_{i}$, $i\in\{1,\dots,k\}$ for each input $x$, using a range of temperature values $\tau=\{0,0.3,0.5,0.7,1.0\}$ to promote output diversity. Each response $y_{i}$ receives a reward $r_{y_{i}}$, and for each pair $(y_{i},y_{j})$, we compute the reward gap $|r_{y_{i}}-r_{y_{j}}|$. Pairs with a gap greater than the threshold $\eta=0.2$ are added to the preference dataset $\mathcal{D}_{p}$. For any such pair where $r_{y_{i}}\geq r_{y_{j}}+\eta$, we assign $y_{win}=y_{i}$, $y_{lose}=y_{j}$, and store the instance $(x,y_{win},y_{lose})\in\mathcal{D}_{p}$. Following Khaki et al. (2024), we train the preference-tuned model $\Phi^{RL}$ using the DPO objective:
$\Phi^{RL}=\arg\max\sum_{(x,y_{win},y_{lose})\in\mathcal{D}_{p}}\log\sigma\left(\beta\log\frac{\Phi^{RL}(y_{win}|x)}{\Phi^{SFT}(y_{win}|x)}-\beta\log\frac{\Phi^{RL}(y_{lose}|x)}{\Phi^{SFT}(y_{lose}|x)}\right)$
@@ 3.4 Training with Reasoning Traces
Finally, inspired by the success of reasoning-augmented models in code and math tasks, and their demonstrated effectiveness in improving CWE detection performance in LLMs (Khare et al., 2023), we propose SFT and DPO methods that incorporate chain-of-thought (CoT) reasoning. To obtain CoT traces that guide the LLM to correct answers, we adopt a rejection sampling approach (``Rejection Sampling SFT'' in Figure 2) for SFT data collection. For each input $x$, we sample $k=5$ responses $y_{i}$, $i\in\{1,\dots,k\}$, from a base untrained CoT-capable LLM (e.g., Qwen3-4B), and compute a reward $r_{y_{i}}$ for each, following the RS-DPO procedure in Figure 2. Instead of forming contrastive pairs, we discard any $y_{i}$ with $r_{y_{i}}<\gamma$, where $\gamma=1$, i.e., CoT–response pairs that are incorrect or improperly formatted. Rewards are applied only to the final response, obtained after parsing the CoT trace, and we also remove cases where the CoT fails to terminate or yield an answer. If no valid $y_{i}$ is found for an input $x$, we skip it. To promote meta-task diversity, we retain at most two valid responses per input: multiple $y_{i}$ only for violation cases and a single $y_{i}$ otherwise. This maintains the 71:29 no-violation-to-violation ratio of Ruff Python SFT data, with the latter more likely to fit within token limits. When excess valid responses exist, we keep the shortest completions, as they typically reflect more concise reasoning (final answers are of similar token length across samples). Following this policy, we collect 52.7k Python training instances from Ruff data, which we use to train the reasoning-enabled base Qwen3-4B with SFT. This yields a CoT-capable SFT model $\Phi_{CoT}^{SFT}$ for Python code quality analysis. We then apply the RS-DPO procedure in Section 3.3 and Figure 2, with the only change being that each $y_{i}$ now includes both the CoT trace and final response.
@@ 4.1 Evaluation Metrics
We evaluate the LLM’s ability to detect idiom violations through two tasks: detection, which assesses whether a given idiom is violated in a code file, and localization, which evaluates whether the model accurately identifies the specific line numbers where the violation occurs. For both tasks, we report precision, recall, and F-score metrics. Detection metrics are calculated at the corpus level for each idiom, treating each as a separate class, while localization metrics are computed at the instance level using set-based precision, recall, and F-score for the ground truth and predicted sets of violating line numbers. To handle potential class imbalance, we use macro-averaging across idioms and exclude NO VIOLATION as a class to penalize models that only predict NO VIOLATIONS FOUND (such models will score zero on all detection metrics). For localization, metrics are averaged only across instances with at least one line of idiom violation in the ground truth. Details of the formal definitions and exact computations of precision, recall, and F-scores for detection and localization are provided in Appendix D.1.
4.2 Generalization on Synthetic Data
To evaluate whether MetaLint training produces adaptive LLMs that handle evolving best practices and novel idioms at test time, we explore transfer settings spanning Python & Java.
Ruff Python Idioms: We construct a 5.3k-instance synthetic test set spanning 50 Ruff idioms, using the data generation procedure from Section 3.1. The data has a 74:26 no-violation-to-violation split, similar to the SFT training set. Idioms are chosen to vary in overlap with training idioms (Figure 3) and fall into three categories:
In domain. 5 idioms identical to those in SFT training, to assess whether MetaLint improves performance on explicitly trained idioms.
@@ 4.2 Generalization on Synthetic Data
PMD and JEP Tree-Sitter Idioms: For Java, we construct two synthetic test sets: 5.1k instances (54:46 split) spanning 269 PMD idioms, and 6.4k instances (50:50 split) spanning 15 JEP idioms (Table 5), flagged via tree-sitter queries. We evaluate in-domain performance by training the base LLM on the corresponding training set (Section 3.2), and also study transfer between PMD and JEP idioms to test adaptation to novel Java idioms. These experiments use Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct to assess the effect of model scale.
4.3 PEP Hard Idiom Benchmark
Benchmark Construction: To test whether MetaLint helps LLMs interpret high-level idiom specifications and generalize to nuanced idioms that linters miss, we construct a benchmark of ``hard idioms'' from 15 PEPs defining semantic or abstract behaviors beyond syntax. We design heuristics per PEP (Table 13, 14 and 15) to detect guideline violations and search the STACK-V2 corpus, prioritizing recall to retrieve broad candidate sets for manual selection. These idioms cannot be reliably detected by simple pattern matching, making them ideal for evaluating model's true understanding versus rote memorization. From the candidates, we handpick 15–20 representative files per PEP, and annotate the precise line ranges (“start” and “end”) of the violated code, providing ground-truth for localization. We add negative examples for each PEP by picking files retrieved for a different PEP and making sure the current PEP is not violated, in order to have a balanced distribution of violation and no-violation cases. The final benchmark contains 536 examples (52% violations, 48% violation-free), enabling evaluation of MetaLint’s generalization from easy to hard Python idioms.
Evaluating Easy-to-Hard Generalization: We use the PEP hard idiom benchmark to test whether training on synthetic data for linter-detectable idioms improves performance on hard idioms. We evaluate the base model, SFT, and DPO-trained models on this benchmark.
Benchmarking on Hard Idioms: We evaluate state-of-the-art open and closed-source code and reasoning LLMs on the PEP hard idiom benchmark, comparing them to MetaLint-trained models. Open-source models include instruction-tuned Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), Qwen3 (Team, 2025), and GPT-oss 20B/120B Agarwal et al. (2025). Closed-source models include GPT-4o (Hurst et al., 2024), o3 mini and o4 mini (OpenAI, 2025a), GPT-4.1 (OpenAI, 2025), and GPT-5 OpenAI (2025b). We select these models for their strong coding and reasoning performance and also evaluate the effects of code-specific pre-training, model scale (3B–120B), and test-time compute for open-source models.
@@ 5 Results
To test whether MetaLint training leads to cross-idiom generalization instead of mere memorization of the training idioms and whether it can produce models that can keep up with evolving code quality standards, we present the transfer performance on the synthetic data for ``easy'' idioms in section 5.1. Then we explore the extent to which MetaLint training achieves easy-to-hard generalization from the synthetic easy idioms to hard, manually curated PEP idioms in section 5.2. Finally, we compare MetaLint trained models against state-of-the-art code and reasoning models on the manually curated hard PEP idioms in section 5.3.
5.1 Generalization on Synthetic Data
Python Ruff Idioms: The performance of Qwen3-4B with and without reasoning and Llama3.2-3B-Instruct when trained on synthetic Ruff idioms and evaluated on the Ruff synthetic test set with varying transfer settings (section 4.2) is shown in Table 1 (full results in Table 18). While Table 1 shows the overall performance, we also analyze the performance broken down by each transfer setting in Table 16. The results show that the SFT stage leads to modest gains in detection and localization performance in most cases (except for a detection recall drop in the case of Llama3.2-3B-Instruct), but the DPO stage leads to huge gains in detection recall, F-score, and all localization metrics at the cost of a slight drop in detection precision. We identify that the drop in precision in the DPO stage is tightly controlled by the fraction of cases with no violations used in the DPO training and explore it in detail in Appendix D.3. Additionally, Table 16 shows that while SFT can lead to slight gains for the transfer settings (near transfer and far transfer), most gains emerge in the DPO stage, especially for non-reasoning models and detection recall. Overall this suggests that SFT can lead to memorization of the training idioms while DPO leads to generalization to novel idioms.
@@ 5.1 Generalization on Synthetic Data
PMD and JEP Tree-Sitter Idioms: To demonstrate the generality of MetaLint training across programming languages and linters, we present results from training on PMD and JEP Tree-Sitter synthetic data in Table 2 (full results in Table 28). Training on PMD shows the same overall pattern as before but with larger recall gains for both SFT and DPO, and notably stronger localization under DPO. For Llama3.1‑8B‑Instruct, SFT initially reduces detection precision, which DPO then recovers; the same precision dip‑and‑recovery appears when transferring PMD→JEP for Llama3.2‑3B‑Instruct. Despite never seeing JEP idioms during training, DPO models achieve strong detection and localization on JEP. In the untrained setting, Llama3.2‑3B‑Instruct (on PMD) and Llama3.1‑3B‑Instruct (on JEP) nearly always output the correct format but predict NO VIOLATIONS FOUND, yielding zero or near‑zero scores because our metrics exclude that class for detection and only score positive cases for localization. Training on JEP yields high in‑domain performance for all metrics with minimal additional benefit from DPO, likely due to JEP’s smaller idiom set (15 vs 269 for PMD) and more precise instructions (Table 5). In the harder JEP→PMD transfer, DPO outperforms SFT, though overall transfer remains weaker than PMD→JEP, reflecting PMD’s broader diversity and more challenging specifications (Appendix C.5).
Overall, MetaLint training consistently yields more adaptable models than the base model, but performance depends on the diversity of training idioms and the gap in instruction quality between training and test data.
@@
5.2 Evaluating Easy-to-Hard Generalizations
To evaluate whether MetaLint training on easy, linter-detectable Ruff idioms improves performance on hard, manually curated PEP idioms, we report results on our PEP hard idiom benchmark (Table 3, full results in Table 19). At the SFT stage, performance declines for Qwen3-4B (with and without CoT) but improves slightly for Llama3.2-3B-Instruct, suggesting that SFT can induce memorization of the training distribution and reduce adaptability. In contrast, DPO yields clear improvements in detection and localization (except detection precision for Llama3.2-3B-Instruct), with statistically significant gains (Appendix E.2). An additional experiment training Qwen3-4B (CoT) directly with RS-DPO, bypassing SFT, resulted in near-zero performance because many generated DPO pairs violated the required output format, which the model inherited. Thus, SFT, despite its drawbacks, is essential for teaching format compliance and setting the stage for DPO to unlock easy-to-hard generalization. Interestingly, the non-CoT model achieves substantially higher detection recall and slightly higher F-score than the CoT variant, despite lower precision. Our analysis attributes the CoT model’s reduced recall to its more conservative interpretation of idiom specifications and to errors such as misinterpretation, overthinking, and skipped lines, as detailed in Appendix E.3.

### v3

@@ 3.2 Instruction Supervised Fine-Tuning
Instruction fine-tuning serves to align the model's outputs with the natural language best practice specifications using the linter outputs as a reference for easy cases. We train the target LLM $\Phi$ on a set of linter-detectable, easy best practice idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$. The input to the model consists of a prompt $p$, which combines a meta-task specification $M_{I}$ for some $I\in I_{\mathcal{L}}$ with a source code file $f\in\mathcal{F}$. The model's output is a list of best practice violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix E.1). In cases where there are no violations ($|V_{f,I}|=0$), the model is expected to output the phrase NO VIOLATIONS FOUND. This formulation explicitly penalizes hallucinated violations of best practices not in the prompt, encouraging the model to flag only violations of the provided best practice and discouraging memorization. We try to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some Python idioms, the final distribution is approximately 70:30 in favor of files with no violations for Python, but roughly 53:47 (PMD) and 50:50 (JEP Tree-Sitter) for the Java data. This results in a total of 53k synthetic training instances spanning 50 idioms (a subset of all the idioms detectable by Ruff) for Python Ruff data and 96.8k instances spanning 269 idioms for Java PMD and 127.3k instances spanning 15 idioms for tree-sitter data, respectively.
@@ 3.3 Verifiable Reward Model and Preference Optimization
We use preference optimization to further improve consistent application of the best practice specification by contrasting accurately localized violations with inaccurate ones, as determined by linter results. We adopt the RS-DPO approach (Khaki et al., 2024), which combines rejection sampling (RS) (Touvron et al., 2023) with Direct Preference Optimization (DPO) (Rafailov et al., 2023) to generate on-policy data from a supervised fine-tuned (SFT) policy model. It samples $k$ outputs per input, computes rewards for them, and constructs contrastive win–loss pairs based on the reward distribution and a threshold $\eta$ (Figure 2). We detail the verifiable linter-based reward model and contrastive pair sampling procedure below.
Reward Model Design: The reward model evaluates model outputs by comparing predicted violations against those flagged by the linter, treating the linter's line numbers (blue circle in ``Verifiable Reward Model'', Figure 1) as ground truth and the model’s predicted lines (yellow circle) as predictions. Reward is computed using set-based precision, recall, and F1-score (visualized via the Venn diagram in the same figure), based on line-level overlap. Since each meta-task $M_{I}$ corresponds to a single best practice idiom $I$, we compute one F1-score (reward) per instance.
Sampling Contrastive Pairs: We begin with an SFT policy model $\Phi^{SFT}$ and sample $k=5$ outputs $y_{i}$, $i\in\{1,\dots,k\}$ for each input $x$, using a range of temperature values $\tau=\{0,0.3,0.5,0.7,1.0\}$ to promote output diversity. Each response $y_{i}$ receives a reward $r_{y_{i}}$, and for each pair $(y_{i},y_{j})$, we compute the reward gap $|r_{y_{i}}-r_{y_{j}}|$. Pairs with a gap greater than the threshold $\eta=0.2$ are added to the preference dataset $\mathcal{D}_{p}$. For any such pair where $r_{y_{i}}\geq r_{y_{j}}+\eta$, we assign $y_{w}=y_{i}$, $y_{l}=y_{j}$, and store the instance $(x,y_{w},y_{l})\in\mathcal{D}_{p}$. By constructing contrastive pairs within the same meta-task, RS-DPO encourages the model to apply a given specification consistently across variations in code structure, rather than across best practices. Following Khaki et al. (2024), we train the preference-tuned model $\Phi^{RL}$ using the DPO objective:
$\Phi^{RL}=\arg\max\sum_{(x,y_{w},y_{l})\in\mathcal{D}_{p}}\log\sigma\Big($ $\beta\log\frac{\Phi^{RL}(y_{w}\mid x)}{\Phi^{SFT}(y_{w}\mid x)}$
$-\beta\log\frac{\Phi^{RL}(y_{l}\mid x)}{\Phi^{SFT}(y_{l}\mid x)}\Big)$
@@ 3.4 Training with Reasoning Traces
Reasoning traces can act as explicit intermediate representations to support transfer across best practice idioms. To obtain CoT traces that guide the LLM to correct answers, we adopt a rejection sampling approach (``Rejection Sampling SFT'' in Figure 2) for SFT data collection. For each input $x$, we sample $k=5$ responses $y_{i}$, $i\in\{1,\dots,k\}$, from a base untrained CoT-capable LLM (e.g., Qwen3-4B), and compute a reward $r_{y_{i}}$ for each, following the RS-DPO procedure in Figure 2. Instead of forming contrastive pairs, we discard any $y_{i}$ with $r_{y_{i}}<\gamma$, where $\gamma=1$, i.e., CoT–response pairs that are incorrect or improperly formatted. Rewards are applied only to the final response, obtained after parsing the CoT trace, and we also remove cases where the CoT fails to terminate or yield an answer. If no valid $y_{i}$ is found for an input $x$, we skip it. To promote meta-task diversity, we retain at most two valid responses per input: multiple $y_{i}$ only for violation cases and a single $y_{i}$ otherwise. This maintains the 71:29 no-violation-to-violation ratio of Ruff Python SFT data, with the latter more likely to fit within token limits. When excess valid responses exist, we keep the shortest completions, as they typically reflect more concise reasoning (final answers are of similar token length across samples). Following this policy, we collect 52.7k Python training instances from Ruff data, which we use to train the reasoning-enabled base Qwen3-4B with SFT. This yields a CoT-capable SFT model $\Phi_{CoT}^{SFT}$ for Python code quality analysis. We then apply the RS-DPO procedure in Section 3.3 and Figure 2, with the only change being that each $y_{i}$ now includes both the CoT trace and final response.
@@ 4.1 Evaluation Metrics
We evaluate the LLM's ability to detect best practice violations through two tasks: detection, which assesses whether a best practice is violated at least once in a code file, and localization, which evaluates whether the model accurately identifies the lines with the problematic idiom. For both tasks, we report precision, recall, and F-score metrics. Detection metrics are calculated at the corpus level for each best practice idiom, treating each as a separate class, while localization metrics are computed at the instance level using set-based precision, recall, and F-score for the ground truth and predicted sets of violating line numbers. To handle potential class imbalance, we use macro-averaging across idioms and exclude NO VIOLATION as a class to penalize models that only predict NO VIOLATIONS FOUND (such models will score zero on all detection metrics). For localization, metrics are averaged only across instances with at least one violation in the ground truth. Details of the formal definitions and exact computations of precision, recall, and F-scores for detection and localization are provided in Appendix F.1.
4.2 Generalization across Easy Best Practices
To evaluate whether MetaLint training encourages transfer to novel easy-to-detect best practices at test time, we explore transfer settings spanning Python & Java.
Ruff Python Idioms: We construct a 5.3k-instance synthetic test set spanning 50 Ruff idioms, using the data generation procedure from Section 3.1. The data has a 74:26 no-violation-to-violation split, similar to the SFT training set. Best practice idioms are chosen to vary in overlap with training idioms (Figure 3) and fall into three categories:
In domain. 5 idioms seen during training, to assess whether MetaLint improves performance on in-domain idioms.
@@ 4.2 Generalization across Easy Best Practices
PMD and JEP Tree-Sitter Idioms: For Java, we construct two synthetic test sets: 5.1k instances (54:46 split) spanning 269 PMD idioms, and 6.4k instances (50:50 split) spanning 15 JEP idioms (Table 5), flagged via tree-sitter queries. We evaluate in-domain performance by training the base LLM on the corresponding training set (Section 3.2), and also study transfer from PMD to JEP and vice-versa. These experiments use Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct to assess the effect of model scale.
4.3 PEP Hard Best Practice Benchmark
To test whether MetaLint trained LLMs exhibit easy-to-hard generalization at test time, we construct a benchmark of hard-to-detect best practices from 15 PEPs that require abstract, context-dependent reasoning. We design heuristics per PEP (Table 15, 16 and 17) to find candidate violations in the STACK-V2 corpus, prioritizing recall to retrieve a broad set for manual selection. These best practices violations cannot be reliably detected by rule-based pattern matching, making them ideal for evaluating a model's contextual reasoning capabilities. From the candidates, we handpick 15–20 representative files per PEP, and annotate the precise line ranges (``start'' and ``end'') of the violated code, providing ground-truth for localization. We add negative examples for each PEP by picking files retrieved for a different PEP and making sure the current PEP is not violated, in order to have a balanced distribution of violation and no-violation cases. The final benchmark contains 536 examples (52% violation, 48% no-violation), enabling evaluation of easy-to-hard generalization by comparing the base model, against the SFT, and DPO-trained models. We also compare MetaLint models trained on Ruff best practices against state-of-the-art open and closed-source code and reasoning LLMs. Open-source models include instruction-tuned Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), Qwen3 (Team, 2025), and GPT-oss 20B/120B (Agarwal et al., 2025). Closed-source models include GPT-4o (Hurst et al., 2024), o3 mini and o4 mini (OpenAI, 2025a), GPT-4.1 (OpenAI, 2025), and GPT-5 (OpenAI, 2025b). We select these models for their strong coding and reasoning performance and also evaluate the effects of code-specific pre-training, model scale (3B–120B), and test-time compute for open-source models.
@@ 5 Results
We evaluate the effectiveness of MetaLint along three dimensions: (1) generalization to unseen easy best practices (Section 5.1); (2) easy-to-hard generalization across best practices (Section 5.2); and (3) whether MetaLint enables a 4B model to achieve performance competitive with state-of-the-art code and reasoning models (Section 5.3).
5.1 Generalization across Easy Best Practices
Python Ruff Idioms: The performance of Qwen3-4B with and without reasoning and Llama3.2-3B-Instruct when trained on synthetic Ruff idioms and evaluated on the Ruff synthetic test set with varying transfer settings (section 4.2) is shown in Table 1 (full results in Table 21). While Table 1 shows the overall performance, we also analyze the performance broken down by each transfer setting in Table 18. The results show that the SFT stage leads to modest gains in detection and localization performance in most cases (except for a detection recall drop for Llama3.2-3B-Instruct), but the DPO stage leads to huge gains in detection recall, F-score, and all localization metrics with a slight drop in detection precision. We identify that the drop in precision in the DPO stage is tightly controlled by the fraction of cases with no violations used in the DPO training and explore it in detail in Appendix F.3. Additionally, Table 18 shows that while SFT can lead to slight gains for the transfer settings, most gains emerge in the DPO stage, especially for non-reasoning models and detection recall. Overall this suggests that SFT is prone to memorizing best practices used for training, while DPO generalizes to novel best practices.
PMD and JEP Tree-Sitter Idioms: To evaluate the generality of MetaLint training across programming languages and linters, we present results from training on PMD and JEP Tree-Sitter synthetic data in Table 2 (full results in Table 31). Training on PMD shows the same overall pattern as before but with larger recall gains for both SFT and DPO, and notably stronger localization under DPO. For Llama3.1‑8B‑Instruct, SFT initially reduces detection precision, which DPO then recovers; the same precision dip‑and‑recovery appears when transferring PMD→JEP for Llama3.2‑3B‑Instruct. Despite never seeing JEP idioms during training, DPO models achieve strong detection and localization on JEP. In the untrained setting, Llama3.2‑3B‑Instruct (on PMD) and Llama3.1‑3B‑Instruct (on JEP) nearly always output the correct format but predict NO VIOLATIONS FOUND, yielding zero or near‑zero scores, since our metrics exclude that class for detection and only score positive cases for localization. Training on JEP yields high in‑domain performance for all metrics with minimal additional benefit from DPO, likely due to JEP’s smaller idiom set (15 vs 269 for PMD) and more precise instructions (Table 5). In the harder JEP→PMD transfer, DPO outperforms SFT, though overall transfer remains weaker than PMD→JEP, reflecting PMD's broader diversity and more challenging specifications (Appendix E.5).
Overall, MetaLint outperforms the base model on novel best practice violations, though gains depend on the diversity of training idioms and the gap in instruction quality between training and test data.
@@ 5.1 Generalization across Easy Best Practices
5.2 Evaluating Easy-to-Hard Generalization
To evaluate whether MetaLint training on easy to detect Ruff best practices improves performance on hard, PEP best practices, we report results on our PEP hard best practice benchmark (Table 3, full results in Table 22). At the SFT stage, performance declines for Qwen3-4B (with and without CoT) but improves slightly for Llama3.2-3B-Instruct, suggesting that SFT is prone to memorization of the training distribution. In contrast, DPO yields clear improvements in detection and localization (except detection precision for Llama3.2-3B-Instruct), with statistically significant gains (Appendix G.2). An additional experiment training Qwen3-4B (CoT) directly with RS-DPO, bypassing SFT, resulted in near-zero performance because many generated DPO pairs violated the required output format, which the model inherited. Thus, SFT, despite its drawbacks, is essential for teaching format compliance and setting the stage for DPO to encourage easy-to-hard generalization. Interestingly, the non-CoT model achieves substantially higher detection recall and slightly higher F-score than the CoT variant, despite lower precision. Our analysis attributes the CoT model's reduced recall to its more conservative interpretation of idiom specifications and to errors such as misinterpretation, overthinking, and skipped lines, as detailed in Appendix G.3. Finally we analyze the impact of the illustrative examples in the prompt in Table 19.
Table 3: Easy-to-Hard Generalization: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3). Best score across the compared training setups per model are bolded.